OpenStudy (anonymous):

The tax on the Reids apartment is $1200 a year. Last year they sold it. They owed only 3 months of the year. How much Tax they pay on the apartment for that year?

OpenStudy (gw2011):

Since they owe 3 months, this becomes 3/12=1/4 (1200)(1/4)=$300 tax